A soccer ball is kicked horizontally off a 22.0-meter high hill and lands a distance of 35.0 meters from the edge of the hill. Determine the initial horizontal velocity of the soccer ball. 




1
Expert's answer
2021-10-31T18:11:36-0400
"0.5(9.8)t^2=22\\t=2.12\ s\\v=\frac{35}{2.12}=16.5\frac{m}{s}"